• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 15:56:29 on October 28, 1998

    Entry number 3959

    keyword=Beam Time Accounting
    Hall A

    10/28/1998 0800-1600
    Experiment #: E94-010
    Shift Leader: Deur A.
    Shift Workers: J.P. Chen, Zhang B.

    08000900100011001200130014001500Total Time
    Acceptable Beam
    in Use
    22 46 53 18 5 1 5 1 4.48333333333333
    Beam Available
    but Not in Use
    0 0 7 27 55 0 55 0 2.4
    Beam Not Available
    or Unacceptable
    38 14 0 15 0 0 0 0 1.11666666666667
    Accelerator Con-
    figuration Chage
    0 0 0 0 0 0 0 0 0
    Experiment
    Ready
    1 1 53 33 1 1 1 1 7.43333333333333
    Planned Config-
    uration Change
    0 0 7 27 0 0 0 0 0.566666666666667
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 27 0 0 0 0 0.45
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 7 0 0 0 0 0 0.116666666666667
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: 8:38 Beam down: arc power supply tripped twice MCC called a specialist. During the beam down: - NMR measurement without laser - NMR measurement with laser --> taget polarization 31 % - Pedestal run 9:46 beam back 10:53 last run in |_ configuration, E=3.4 GeV, Ep=2.199 GeV, theta=15 degree Tot # of events 17.6 M beam secured during momentum changed 11:27 beam restored. Begin // runs @ E=3.4 GeV, Ep=2.030 GeV, 15.5 degree 11:45 Beam down (PSS fault in the injector) 12:55 beam back (PSS problem not fixed but the need beam to troubleshoot it) 14:00 coda crashed 14:55 coda back.

    Data Entry by: Deur A.